OpenStudy (kendricklamar2014):

Simplify each expression 9! A. 45 B. 362,880 C. 265,456 D. 125,936

OpenStudy (praxer):

9!= 9*8*7*6*5*4*3*2*1= 362,880

OpenStudy (anonymous):

9*8*7*6*5*3*2*1=?

OpenStudy (anonymous):

B

OpenStudy (praxer):

option, B is correct.... :)

OpenStudy (kendricklamar2014):

Thanks

OpenStudy (anonymous):

np love! :)

OpenStudy (kendricklamar2014):

haha :)

OpenStudy (kendricklamar2014):

Can you help me with one more?

OpenStudy (praxer):

$$\huge Welcome...\ddot\smile$$

OpenStudy (kendricklamar2014):

25P3

OpenStudy (praxer):

it is 25!/(25-9)!

OpenStudy (kendricklamar2014):

No this is another question

OpenStudy (kendricklamar2014):

Answer choices: A. 12,755 B. 14,290 C. 13,800 D. 18,250

OpenStudy (praxer):

oh sorry I made a mistake it should be 25!/(25-3)! =25!/22! =25*24*23*22!/22! =25*24*23 =13,800. Option B. is correct..... :)

OpenStudy (praxer):

$$\huge \color {red}{option \ C. is \ correct}$$

OpenStudy (kendricklamar2014):

How do you make the letter big?

OpenStudy (anonymous):

C

OpenStudy (praxer):

http://openstudy.com/study#/updates/52e12f56e4b0942cc9de719e This will help for all the latex..... :)